chocolateluva77:

can smo pls answer asap? algebraically solve the system {-3x + 4y = 7 6x - 8y = 14 pls show how u got the steps to the answer. ty.

Sailor:

Are there 3 separate equations or one equation?

chocolateluva77:

@Sailor it's 2 equations. find x n y

Sailor:

Ok ok

chocolateluva77:

@Sailor do yk how to do it?

surjithayer:

-3x+4y=7 multiply by -2 6x-8y=-14 and 6x-8y=14 so -14=14 which is impossible Hence it has no solution.

Sailor:

@surjithayer wrote:
-3x+4y=7 multiply by -2 6x-8y=-14 and 6x-8y=14 so -14=14 which is impossible Hence it has no solution.
:0 your smart

chocolateluva77:

@surjithayer ty

chocolateluva77:

@Sailor fr i think he helped me on another question too

surjithayer:

yw